# Distilling Machine Learning’s Added Value: Pareto Fronts in Atmospheric Applications

### v1

@@ Abstract
While the added value of machine learning for weather and climate applications is measurable, explaining it remains challenging, especially for large deep learning models. Inspired by climate model hierarchies, we propose that a full hierarchy of Pareto-optimal models, defined within an appropriately determined error-complexity plane, can guide model development and help understand the models’ added value. We demonstrate the use of Pareto fronts in atmospheric physics through three sample applications, with hierarchies ranging from semi-empirical models with minimal tunable parameters (simplest) to deep learning algorithms (most complex). First, in cloud cover parameterization, we find that neural networks identify nonlinear relationships between cloud cover and its thermodynamic environment, and assimilate previously neglected features such as vertical gradients in relative humidity that improve the representation of low cloud cover. This added value is condensed into a ten-parameter equation that rivals the performance of deep learning models. Second, we establish a machine learning model hierarchy for emulating shortwave radiative transfer, distilling the importance of bidirectional vertical connectivity for accurately representing absorption and scattering, especially for multiple cloud layers. Third, we emphasize the importance of convective organization information when modeling the relationship between tropical precipitation and its surrounding environment. We discuss the added value of temporal memory when high-resolution spatial information is unavailable, with implications for precipitation parameterization. Therefore, by comparing data-driven models directly with existing schemes using Pareto optimality, we promote process understanding by hierarchically unveiling system complexity, with the hope of improving the trustworthiness of machine learning models in atmospheric applications.
With machine learning permeating the geosciences, it becomes urgent to distinguish incremental progress from enduring knowledge discovery. We show that Pareto-optimal hierarchies transparently distill the added value of new algorithms using three atmospheric physics applications, providing a timely complement to post-hoc explainable artificial intelligence tools.
@@ 1 Introduction
Have recent advancements in machine learning (ML) led to discoveries in atmospheric science? The added value of ML for weather and climate applications is measurable, but it often remains challenging to understand. Taking advancements in data-driven, medium-range weather forecasting (Ben-Bouallegue et al., 2023) as an example, increasing reliance on complex architectures makes state-of-the-art models difficult to interpret.
@@ 2 Pareto-Optimal Model Hierarchies
In this section, we define Pareto optimality before discussing the definition of error and complexity.
@@ 2.1 Pareto Optimality
Intuitively, when we select a model from the Pareto front, any attempt to switch to a different model would mean sacrificing the quality of at least one evaluation metric. Conversely, a model that can be replaced without compromising any evaluation metrics is described as Pareto-dominated. In practice, we often seek to balance evaluation metrics measuring error and complexity, which we discuss in the next subsections.
@@ 2.2 Error
We emphasize the importance of holistic evaluation, which employs several error metrics with different behaviors: Traditional regression or classification metrics, distributional distances, spectral metrics, probabilistic scoring rules, reliability diagrams (Haynes et al., 2023), causal evaluation (Nowack et al., 2020), etc. To facilitate the use of Pareto-optimal hierarchies, we recommend prioritizing proper scores, whose expectation is optimal if and only if the model represents the true target distribution (Bröcker, 2009). For simplicity’s sake, we will employ mean squared error (MSE) as our primary error metric for our study’s applications. We make this choice because MSE is a proper score for deterministic predictions that can be efficiently optimized, while recognizing MSE’s inherent limitations for non-normally distributed targets.
@@ 2.3 Complexity
To our knowledge, there are no universally accepted metrics for quantifying the complexity of data-driven models within the geosciences. In statistical learning, various complexity metrics, such as Rademacher complexity (e.g., Bartlett et al., 2005), rely on dataset characteristics, whereas others, like the VC dimension (Vapnik and Chervonenkis, 2015), solely depend on algorithmic attributes. Here, we predominantly focus on two metrics that can be readily calculated from model attributes: the number of trainable parameters () and the number of features (). We choose these metrics due to their simplicity and versatility across the broad spectrum of models considered in our study. As we will confirm empirically in Sec. 4, and can be used as (very) approximate proxies for generalizability, as models with fewer trainable parameters and features with more stable distributions tend to result in superior generalizability. We defer the exploration of additional complexity metrics, such as the number of floating point operations (FLOP), to future work.
@@ 3 The Distillable Value of Machine Learning
In geoscientific applications, we often work with features that are functions of space, discretized in spatial locations , and of time, discretized in timesteps . We will consider a model predicting a target vector from a multi-variate spatiotemporal field such that . We define a model as having added value w.r.t. a set of evaluation metrics and a simple baseline model if the following is true: When applied to representative out-of-sample data, is Pareto-optimal while the baseline model is not. Taking the data’s spatiotemporal structure into account, we organize a model’s added value into four mutually non-exclusive categories: functional representation, feature assimilation, spatial connectivity, and temporal connectivity (see Fig. 1).
@@ 3.1 Functional Representation
A fundamental aspect a model can improve is the functional representation of the observed relationship between a set of baseline features —fixed features used to benchmark performance—and the target . We deem to improve the functional representation of a baseline w.r.t. if and only if:
This improvement may stem from algorithms leading to better fits (e.g., gradient boosting instead of decision trees), improved optimization (e.g., the Adam optimizer and its variants instead of traditional stochastic gradient descent; Kingma and Ba, 2014), improved parsimony (e.g., by decreasing the number of trainable parameters via hyperparameter tuning) or enforced constraints (e.g., positive concentrations and precipitation). Improvements in functional representation are readily visualized via partial dependence plots or their variants (marginal plots to avoid unlikely data instances, accumulated local effects to account for feature correlation, etc.; Molnar, 2020). Note that Eq. 2 also captures improvements in probabilistic modeling by generalizing to a stochastic mapping and including probabilistic scores (Gneiting and Raftery, 2007; Haynes et al., 2023) in . From an atmospheric science perspective, improving functional representation helps identify nonlinear regimes, identify model extremes, and describe how sensitive the prediction is to different features. However, it can be challenging to faithfully describe these sensitivities if key features are missing from the baseline set of features .
@@ 4 Atmospheric Physics Application Cases
Building upon the four types of added value outlined in the previous section, this section demonstrates that Pareto optimality guides model development and improves process understanding through three realistic, atmospheric modeling case studies. Each case includes machine learning prototypes with demonstrated performance from previous studies, along with Pareto-optimal models newly trained for this study.
@@ 4.1.1 Motivation
Storm-resolving simulations with grid spacing below km reduce uncertainty related to the interaction between storms and planetary-scale dynamics by explicitly simulating deep convection (Stevens et al., 2020). However, their large computational cost prohibits their routine use for ensemble projections (Schneider et al., 2017). Machine learning can learn the storm-scale behavior of clouds from short, storm-resolving simulations, potentially improving coarser Earth system models through data-driven parameterizations (Gentine et al., 2021).
In this section, we aim to understand the improvement gained from the higher-fidelity representation of storms and clouds. As illustrated in Fig. 2, we demonstrate that this knowledge can be symbolically distilled into an analytic equation that rivals the performance of deep learning.
@@ 4.1.2 Setup
Once coarse-grained and pre-processed, we aim to map the environmental conditions on vertical levels indexed by the background terrain-following height grid , to the cloud cover on the same vertical levels. The variables include the horizontal wind speed [m/s], specific humidity [kg/kg], liquid water mixing ratio [kg/kg], ice mixing ratio [kg/kg], temperature [K], pressure [Pa], and relative humidity . Except for the “non-local NN” in Sec. 4.4.1.4.1.4, we simplify the mapping to a “vertically quasi-local” one where cloud cover at a given level depends only on the atmospheric variables at the same level and their first and second-order derivative with respect to . also includes geometric height [m] and surface variables: land fraction [%] and surface pressure [Pa]. In summary, we approximate some mapping:
using a hierarchy of machine learning models. In the following subsections, we will show that Pareto-optimal hierarchies not only facilitate data-driven model development, allowing for the comparison of simple baselines with neural networks, but also promote process understanding when tracing the Pareto front.
4.1.3 Existing Baselines and Polynomial Regression
@@
To account for marine stratocumulus (low) clouds (Mauritsen et al., 2019), we use two different sets of 4 parameters for land and sea using a land fraction threshold of 0.5. The Sundqvist scheme is parsimonious with only 8 trainable parameters, but it performs poorly against high-resolution data, with MSE values as large as despite having been re-tuned to our training set.
@@
i.e., that the absence of condensates implies no clouds, further improves the results. Note that this constraint can also be derived from data using a binary decision tree. Accounting for this constraint, polynomial models achieve MSEs less than half of the Sundqvist scheme’s (see Fig 2’s lower error, higher complexity polynomial models). Focusing on the Pareto-optimal, multilinear and quadratic model outputs:
with parameters listed in appendix A. Analyzing this first Pareto front unveils the role of each added feature in better representing cloud cover. First, it highlights the role of cloud condensates, both liquid and ice, in accurately describing cloud cover. Second, among all possible vertical gradients, Pareto-optimal models select the negative gradient of relative humidity to depict the role of inversions in increasing cloud cover, which is particularly important for low clouds, including marine stratocumulus clouds. However, upon closer inspection, these contributions are not always physically consistent: while the multilinear model correctly captures the increase of cloud cover with ice, it incorrectly decreases cloud cover when there is humidity above. Conversely, the quadratic model accurately captures the inversion role when temperature and humidity are fixed to their training set mean but wrongly assumes that inversions’ effect on cloudiness depends on humidity and temperature, and the model’s decrease in cloudiness with increasing mixed-phase cloud condensates suggests that its term is not an interpretable sensitivity to liquid and ice but rather a bias correction term relying on these variables.
@@
with in appendix A, suggests that the dependence of cloud cover on condensates is fundamentally non-linear. This justifies the use of deep learning to quickly explore which features can be combined to improve the non-linear, functional representation of cloud cover.
4.1.4 Deep Learning
We start with the baseline model from Grundner et al. (2022), which uses a 3-layer-deep, 64-neuron-wide multi-layer perceptron with batch normalization after the second hidden layer. Hyperparameters were optimized using the SHERPA Python library (Hertel et al., 2020). This NN estimates the target cloud cover with high fidelity (MSE=), but at the cost of increased complexity, as it has a total of 9345 trainable parameters. Note that the models from Grundner et al. (2022) were not designed to minimize the number of trainable parameters, so we do not overly focus on this complexity metric. The MSE can be further lowered to with vertically non-local NNs, which map the entire atmospheric column of inputs to the entire column of outputs without inductive bias. However, this small error gain is deemed insufficient given the increased complexity cost.
@@ 4.1.4 Deep Learning
Unlike features selected by polynomial models, these can be non-linearly combined to yield high-quality predictions, as shown in the right panels of Fig. 2. First, NNs improve the functional representation of cloud cover by accurately modeling the sharp increase in cloud cover above a temperature-dependent relative humidity threshold – a highly nonlinear, bivariate behavior that simple schemes struggle to capture. Formally:
@@ 4.1.5 Symbolic Distillation and Equation Discovery
These are in fact emerging properties of the subgrid distributions of liquid and ice (Grundner et al., 2024): As large values of cloud ice are rarely observed, larger spatial averages of cloud ice at coarse resolution means that many more high-resolution pixels contain low values of cloud ice compared to the liquid case, resulting in higher cloudiness for a given spatially-averaged condensate value. By assuming an exponential distribution for the subgrid liquid and ice content, we can even interpret and as the rate of the respective exponential distributions. This allows us to hypothesize that the nonlinear relationship between condensates and cloud cover is not scale-invariant and requires separate treatments of liquid and ice, with implications for the interaction between microphysical processes and the radiative budget. While analyzing the feature importance of liquid and ice in neural networks could have suggested this difference, it would have been difficult to fully explain it and bridge spatial scales without distillation, confirming the importance of only treating deep learning as a first and not final step towards knowledge discovery. We now turn to a higher-dimensional problem in which spatial connectivity has to be considered: radiative transfer.
@@ 4.2.1 Motivation
In this section, we examine how ML architectural designs impact the reliability of shortwave radiative transfer (covering solar radiation and wavelengths of 0.23-5.85 m). As shown in Fig. 3, architectures that closely mimic the vertical bidirectionality of radiative transfer are Pareto-optimal, rivaling the performance of deep learning models with ten times more trainable parameters.
@@ 4.2.2 Setup
4.2.3 Machine Learning Model Hierarchy
We deploy an ML model hierarchy using the same features to isolate the added value of spatial connectivity along the vertical coordinate . For all models, cropping and zero-padding layers set the top-of-atmosphere (TOA; 78 km above ground level) heating rate to zero, consistent with the RRTM model being emulated. Readers interested in the exact ML model architecture are referred to Lagerquist et al. (2022) and the accompanying code (https://zenodo.org/records/12557544).

### v2

@@ Abstract
The added value of machine learning for weather and climate applications is measurable through performance metrics, but explaining it remains challenging, particularly for large deep learning models. Inspired by climate model hierarchies, we propose that a full hierarchy of Pareto-optimal models, defined within an appropriately determined error-complexity plane, can guide model development and help understand the models’ added value. We demonstrate the use of Pareto fronts in atmospheric physics through three sample applications, with hierarchies ranging from semi-empirical models with minimal parameters (simplest) to deep learning algorithms (most complex). First, in cloud cover parameterization, we find that neural networks identify nonlinear relationships between cloud cover and its thermodynamic environment, and assimilate previously neglected features such as vertical gradients in relative humidity that improve the representation of low cloud cover. This added value is condensed into a ten-parameter equation that rivals deep learning models. Second, we establish a machine learning model hierarchy for emulating shortwave radiative transfer, distilling the importance of bidirectional vertical connectivity for accurately representing absorption and scattering, especially for multiple cloud layers. Third, we emphasize the importance of convective organization information when modeling the relationship between tropical precipitation and its surrounding environment. We discuss the added value of temporal memory when high-resolution spatial information is unavailable, with implications for precipitation parameterization. Therefore, by comparing data-driven models directly with existing schemes using Pareto optimality, we promote process understanding by hierarchically unveiling system complexity, with the hope of improving the trustworthiness of machine learning models in atmospheric applications.
As machine learning permeates the geosciences, it becomes urgent to identify improvements in process representation and prediction that challenge our understanding of the underlying system. We show that Pareto-optimal hierarchies transparently distill the added value of new algorithms in three atmospheric physics applications, providing a timely complement to post-hoc explainable artificial intelligence tools.
@@ 1 Introduction
Will recent advancements in machine learning (ML) lead to enduring new knowledge in atmospheric science? While the added value of ML for weather and climate applications can be measured using performance metrics, it often remains challenging to understand. Taking advancements in data-driven, medium-range weather forecasting (Ben-Bouallegue et al., 2023) as an example, increasing reliance on complex architectures makes state-of-the-art models difficult to interpret.
@@ 2.1 Pareto Optimality
Intuitively, when we select a model from the Pareto front, any attempt to switch to a different model would mean sacrificing the quality of at least one evaluation metric. Conversely, a model that can be replaced without compromising any evaluation metrics is described as Pareto-dominated. Henceforth, we derive Pareto fronts empirically from the available data and the subset of models considered. These empirical Pareto fronts are denoted , acknowledging that the “true Pareto front” is generally intractable. In practice, we often seek to balance evaluation metrics measuring error and complexity, which we discuss in the next subsections.
@@ 2.2 Error
We emphasize the importance of holistic evaluation, which employs several error metrics with different behaviors: Traditional regression or classification metrics, distributional distances, spectral metrics, probabilistic scoring rules, reliability diagrams (Haynes et al., 2023), causal evaluation (Nowack et al., 2020), etc. To facilitate the use of Pareto-optimal hierarchies, we recommend prioritizing proper scores, whose expectation is optimal if and only if the model represents the true target distribution (Bröcker, 2009). For simplicity’s sake, we will employ mean squared error (MSE) and its square root (RMSE) as our primary error metric for our study’s applications. We make this choice because MSE is a proper score for deterministic predictions that can be efficiently optimized, while recognizing MSE’s inherent limitations for non-normally distributed targets.
@@ 2.3 Complexity
To our knowledge, there are no universally accepted metrics for quantifying the complexity of data-driven models within the geosciences. In statistical learning, various complexity metrics, such as Rademacher complexity (e.g., Bartlett et al., 2005), rely on dataset characteristics, whereas others, like the VC dimension (Vapnik and Chervonenkis, 2015), solely depend on algorithmic attributes. Here, we predominantly focus on two metrics that can be readily calculated from model attributes: the number of trainable parameters () and the number of features (). We choose these metrics due to their simplicity and versatility across the broad spectrum of models considered in our study (Balaji et al., 2017). As we will confirm empirically in Sec. 4, and can be used as (very) approximate proxies for generalizability, as models with fewer trainable parameters and features with more stable distributions tend to result in superior generalizability. We defer the exploration of additional complexity metrics, such as the number of floating point operations (FLOP) or Multiply-Accumulate Operations (MACs), to future work.
@@ 3 The Distillable Value of Machine Learning
In geoscientific applications, we often work with features that are functions of space, discretized in spatial locations , and of time, discretized in timesteps . We consider a model predicting a target vector from a multi-variate spatiotemporal field such that:
We define a model as having added value w.r.t. a set of evaluation metrics and a baseline model if the following holds: When applied to representative out-of-sample data, is Pareto-optimal while the baseline model is not. Note that improving the Pareto front is necessary but not sufficient: a model predicting zero everywhere does not generally add value compared to other models solely because it is simpler.
Equation 2 suggests four degrees of freedom for a model’s added value: (1) improving the function without changing the features (functional representation), (2) improving the model through (feature assimilation), (3) improving the model through (spatial connectivity), and (4) improving the model through (temporal connectivity). We depict these four non-mutually-exclusive categories in Fig. 1 and rigorously define them below.
@@ 3.1 Functional Representation
A fundamental aspect a model can improve is the functional representation of the observed relationship between a set of baseline features —fixed features used to benchmark performance—and the target . We deem to improve the functional representation of a baseline w.r.t. and if and only if:
This improvement may stem from algorithms leading to better fits (e.g., gradient boosting instead of decision trees), improved optimization (e.g., the Adam optimizer and its variants instead of traditional stochastic gradient descent; Kingma and Ba, 2014), improved parsimony (e.g., by decreasing the number of trainable parameters via hyperparameter tuning) or enforced constraints (e.g., positive concentrations and precipitation). Improvements in functional representation are readily visualized via partial dependence plots or their variants (marginal plots to avoid unlikely data instances, accumulated local effects to account for feature correlation, etc.; Molnar, 2020). Note that Eq. 3 also captures improvements in probabilistic modeling by generalizing to a stochastic mapping and including probabilistic scores (Gneiting and Raftery, 2007; Haynes et al., 2023) in . From an atmospheric science perspective, improving functional representation helps identify nonlinear regimes, model extremes, and describe how sensitive the prediction is to different features. However, it can be challenging to faithfully describe these sensitivities if key features are missing from the baseline set of features .
@@ 4 Atmospheric Physics Application Cases
This section demonstrates that Pareto optimality guides model development and improves process understanding through three realistic, atmospheric modeling case studies. Each case includes machine learning prototypes with demonstrated performance from previous studies, along with Pareto-optimal models newly trained for this study. The first case study emphasizes functional representation and feature assimilation, the second focuses on spatial connectivity, and the last compares spatial and temporal connectivity.
@@ 4.1.1 Motivation
Although “storm-resolving” simulations with grid spacing below km do not explicitly resolve clouds and their associated microphysical processes (Morrison et al., 2020), they significantly reduce uncertainty in the interaction between storms and planetary-scale dynamics by explicitly simulating deep convection (Stevens et al., 2020). However, their large computational cost prohibits their routine use for ensemble projections (Schneider et al., 2017). Machine learning can learn the storm-scale behavior of clouds from short, storm-resolving simulations, potentially improving coarser Earth system models through data-driven parameterizations (Gentine et al., 2021).
This case study aims to understand the improvement gained from the higher-fidelity representation of storms and clouds. As illustrated in Fig. 2, we demonstrate that this knowledge can be symbolically distilled into an analytic equation that rivals the performance of deep learning.
@@ 4.1.2 Setup
Once coarse-grained and pre-processed, we aim to map the environmental conditions on vertical levels indexed by the background terrain-following height grid , to the cloud cover on the same vertical levels. The variables include the horizontal wind speed [m/s], specific humidity [kg/kg], liquid water mixing ratio [kg/kg], ice mixing ratio [kg/kg], temperature [K], pressure [Pa], and relative humidity . Except for the “non-local NN” in the next sub-section, we simplify the mapping to a “vertically quasi-local” one where cloud cover at a given level depends only on the atmospheric variables at the same level and their first and second-order derivative with respect to . also includes geometric height [m] and surface variables: land fraction [%] and surface pressure [Pa]. In summary, we approximate some mapping:
using a hierarchy of machine learning models. In the following subsections, we will show that Pareto-optimal hierarchies not only facilitate data-driven model development, allowing for the comparison of simple baselines with neural networks, but also promote process understanding.
4.1.3 Model Hierarchy
@@
To account for marine stratocumulus (low) clouds (Mauritsen et al., 2019), we use two different sets of 4 parameters for land and sea using a land fraction threshold of 0.5. The Sundqvist scheme is parsimonious with only 8 trainable parameters, but it performs poorly against high-resolution data, with RMSE values as large as despite having been re-tuned to our training set.
@@
i.e., that the absence of condensates implies no clouds, further improves the results. Note that this constraint can also be derived from data using a binary decision tree. Accounting for this constraint, polynomial models achieve MSEs less than half of the Sundqvist scheme’s (see Fig 2’s lower error, higher complexity polynomial models). Focusing on the Pareto-optimal, four-feature linear and quadratic model outputs:
with parameters listed in appendix A. Analyzing this first Pareto front unveils the role of each added feature in better representing cloud cover. First, it highlights the role of cloud condensates, both liquid and ice, in accurately describing cloud cover. Second, among all possible vertical gradients, Pareto-optimal models select the negative gradient of relative humidity to depict the role of inversions in increasing cloud cover, which is particularly important for low clouds, including marine stratocumulus clouds. However, upon closer inspection, these contributions are not always physically consistent: while the four-feature linear model correctly captures the increase of cloud cover with ice, it incorrectly decreases cloud cover when there is humidity above. Conversely, the quadratic model accurately captures the inversion role when temperature and humidity are fixed to their training set mean but wrongly assumes that inversions’ effect on cloudiness depends on humidity and temperature, and the model’s decrease in cloudiness with increasing mixed-phase cloud condensates suggests that its term is not an interpretable sensitivity to liquid and ice but rather a bias correction term relying on these variables.
@@
with in appendix A, suggests that the dependence of cloud cover on condensates is fundamentally nonlinear. This justifies the use of deep learning to quickly explore which features can be combined to improve the nonlinear, functional representation of cloud cover.
We start with the baseline neural network from Grundner et al. (2022): a 3-layer-deep, 64-neuron-wide multi-layer perceptron with batch normalization after the second hidden layer. Hyperparameters were optimized using the SHERPA Python library (Hertel et al., 2020). This NN estimates the target cloud cover with high fidelity (RMSE=), but at the cost of increased complexity, as it has a total of 9345 trainable parameters. Note that the models from Grundner et al. (2022) were not designed to minimize the number of trainable parameters, so we do not overly focus on this complexity metric. The RMSE can be further lowered to with vertically non-local NNs, which map the entire atmospheric column of inputs to the entire column of outputs without inductive bias. However, this small error gain is deemed insufficient given the increased complexity cost.
@@
Unlike features selected by polynomial models, these can be nonlinearly combined to yield high-quality predictions, as shown in the right panels of Fig. 2. First, NNs improve the functional representation of cloud cover by accurately modeling the sharp increase in cloud cover above a temperature-dependent relative humidity threshold – a highly nonlinear, bivariate behavior that simple schemes struggle to capture. Formally:
@@ 4.1.4 Symbolic Distillation and Equation Discovery
These are in fact emerging properties of the subgrid distributions of liquid and ice (Grundner et al., 2024): As large values of cloud ice are rarely observed, larger spatial averages of cloud ice at coarse resolution means that many more high-resolution pixels contain low values of cloud ice compared to the liquid case, resulting in higher cloudiness for a given spatially-averaged condensate value. By assuming an exponential distribution for the subgrid liquid and ice content, we can even interpret and as the rate of the respective exponential distributions. This allows us to hypothesize that the nonlinear relationship between condensates and cloud cover is not scale-invariant and requires separate treatments of liquid and ice, with implications for the interaction between microphysical processes and the radiative budget. While analyzing the feature importance of liquid and ice in neural networks could have suggested this difference, it would have been difficult to fully explain it and bridge spatial scales without distillation, confirming the importance of only treating deep learning as a first and not final step towards knowledge discovery.
@@ 4.2.1 Motivation
This case study examines how ML architectural designs impact the reliability of shortwave radiative transfer (covering solar radiation and wavelengths of 0.23-5.85 m). As shown in Fig. 3, physics-informed architectures that closely mimic the vertical bidirectionality of radiative transfer are Pareto-optimal, rivaling the performance of deep learning models with ten times more trainable parameters.
@@ 4.2.2 Setup
4.2.3 Model Hierarchy
We deploy an ML model hierarchy using the same features to isolate the added value of spatial connectivity along the vertical coordinate . For all models, cropping and zero-padding layers set the top-of-atmosphere (TOA; 78 km above ground level) heating rate to zero, consistent with the RRTM model being emulated. Readers interested in the exact ML model architecture are referred to Lagerquist et al. (2022).